you have a normal population of scores with u=60 and o=10. we obtained a random sample of n=40. what is the probability that the sample mean will be less than 54?
Answer #1

Solution :

Given that ,

mean = = 60

standard deviation = = 10

n = 40

= 60

=  / n = 10 / 40=1.58

P( < 54) = P[( - ) / < (54-60) /1.58 ]

= P(z <-3.80 )

Using z table  

probability= 0.0001